OpenStudy (anonymous):

for the curve x^3 + y^2 = x + y passing through (-2,3) use the tangent line at x = 2 to approximate the value of y at x = -1.5

OpenStudy (anonymous):

implicit diff?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

start with \[3x^2+3y^2y'=1+y'\]

OpenStudy (anonymous):

wait... where did you get the 3y'y'?

OpenStudy (anonymous):

it is not \(3y'y'\) is it \(3y^2y'\)

OpenStudy (anonymous):

right but why didn't you put 2yy' since its squared?

OpenStudy (anonymous):

oh doh, because i read it wrong sorry it is in fact what you said \[3x^2+2yy'=1+y'\]

OpenStudy (anonymous):

oh ok:)

OpenStudy (anonymous):

so.... how do we approach it next?

OpenStudy (anonymous):

i guess you need the slope of the line \(y'\) at the point \((-2,3)\)

OpenStudy (anonymous):

now that i look at the problem i can't say it makes much sense that slope will have nothing to do with the curve at \(x=2\) are you sure it is not \(x=-2\)?

OpenStudy (anonymous):

oh no you're right! it is -2! sorry!

OpenStudy (anonymous):

whew so now you have a choice you can solve \[3x^2+2yy'=1+y'\] for \(y'\) and then put \(x=-2,y=3\) and see what you get

OpenStudy (anonymous):

or you can put \(x=-2,y=3\) at the start and solve \[3\times(-2)^2+2\times 3\times y'=1+y'\] for \(y\) using elementary algebra either way will give \(y'\) at \((-2,3)\)

OpenStudy (anonymous):

so is the answer y' = 11/12? and is that supposed to be the slope or did i completely miss this?

OpenStudy (aum):

\[ 12 + 6y' = 1 + y' \\ 5y' = -11 \\ y' = -\frac{11}{5} \]

OpenStudy (anonymous):

oh jeez i multiplied wrong...

OpenStudy (anonymous):

then find the equation of the line with slope \[-\frac{11}{5}\] through the point \((-2,3)\) leave it in the point slope form

OpenStudy (aum):

\[ y' = \frac{y_2-y_1}{x_2-x_1} = \frac{y_2-3}{-1.5+2} = -\frac{11}{5} \]

OpenStudy (anonymous):

finally replace \(x\) by \(-1.5\)

OpenStudy (aum):

\[(x_1, y_1) = (-2, 3) \\ (x_2,y_2) = (-1.5, y_2) \\ \text{Slope = }\frac{y_2-y_1}{x_2-x_1} = \frac{y_2-3}{-1.5+2} \]

OpenStudy (anonymous):

-4.1 is y correct?

OpenStudy (aum):

\[ \frac{y_2-3}{-1.5+2} = -\frac{11}{5} \\ y_2-3 = -2.2 * 0.5 = -1.1\\ y_2 = 3 - 1.1 = 1.9 \text{ or } \frac{19}{10} \]

OpenStudy (anonymous):

oh haha i subtracted the 3...

OpenStudy (anonymous):

but thank you guys sooo much!!

OpenStudy (aum):

you are welcome.